# Governance Architecture for Autonomous Agent Systems: Threats, Framework, and Engineering Practice

### 07 Mar 2026 v1

@@ 7.4.6. Quantifying the GPT-4o-mini Circularity Bias
Since GPT-4o-mini generated the benchmark and also serves as one of the evaluated judges, its interception rates may be inflated by familiarity with its own output distribution. We bound this bias by comparing GPT-4o-mini’s performance against Qwen3.5-9B, an independently developed model from a different vendor (Alibaba vs. OpenAI) with a distinct training pipeline. On the Chinese dataset, Qwen3.5-9B achieves 96.5% IR versus GPT-4o-mini’s 93.0%; on English, 95.0% versus 95.4%. The fact that an independent model achieves equal or higher interception strongly suggests that the benchmark captures genuine attack patterns rather than GPT-4o-mini-specific artifacts. Moreover, GPT-4o-mini’s lower interception on Chinese (93.0% vs. 96.5%) contradicts the circularity hypothesis: if same-source bias were dominant, GPT-4o-mini should consistently outperform, which it does not. On IR, GPT-4o-mini exceeds Qwen3.5-9B only on English (95.4% vs. 95.0%, +0.4 pp) while underperforming on Chinese (93.0% vs. 96.5%, pp); the inconsistent direction across languages indicates that any same-source IR inflation is small relative to the inter-model variance. On FPR, GPT-4o-mini’s lower rate (3.2–12.3% vs. 23.7–37.7%) could partly reflect familiarity with its own benign outputs; this effect cannot be isolated without regenerating the benchmark with an independent generator. We note that this analysis bounds output-distribution familiarity on IR; a deeper structural bias—whereby GPT-4o-mini may also recognize its own attack-pattern preferences embedded in the generation templates—is harder to bound and remains a limitation. An analogous bias on FPR, whereby GPT-4o-mini may more reliably classify its own benign outputs as safe, cannot be excluded; bounding either effect would require regenerating the benchmark with an independent model (e.g., Claude or Gemini).
@@ 7.6. Generalization to External Benchmark
Table 10 shows that both judges achieve near-perfect interception on InjecAgent (GPT-4o-mini: 100% [95% CI: 96.4–100%]; Qwen3.5-9B: 99% [95% CI: 94.6–100%]), matching or exceeding their performance on our synthetic indirect injection subset. This confirms that the intent verification approach generalizes to independently constructed attack scenarios and is not overfit to our benchmark’s generation process. Note that these results measure judge-level interception on isolated tool calls, not end-to-end agent compromise reduction; validating LGA’s impact on full-pipeline attack success rates remains future work.
@@ 8. Conclusion
This paper argues that engineering effort in AI-augmented systems is shifting from defect remediation toward system governance and boundary design. We formalize three security threat classes in multi-agent systems—Agency Abuse, RAG Data Poisoning, and Malicious Skill Plugins—and propose LGA, a four-layer governance architecture with empirical evaluation across all three.
RQ1 (Intent verification effectiveness). Across four local LLM judges, two NLI baselines, and one cloud judge evaluated on a 1,081-sample bilingual benchmark, all five LLM-based judges achieve at least 93% overall interception on TC1/TC2, while lightweight NLI baselines remain below 10%. Qwen2.5-14B achieves the best local balance (98.2–98.5% IR, 9.7–20.1% FPR), approaching GPT-4o-mini (93.0–95.4% IR, 3.2–12.3% FPR). On TC3 (malicious skill plugins), interception among judges with meaningful precision–recall balance drops to 75–94%; judges at the high end of the raw range (e.g., Llama-3.1-8B ZH at 99.2%) achieve this via conservative blocking bias (87.5% FPR on benign plugin calls; see Sect. 7.4.3), confirming that permission-boundary violations are harder to detect from intent alignment alone and motivating complementary enforcement at Layers 1 and 3 (sandbox isolation and schema validation). Despite this TC3 limitation, generalization to the external InjecAgent benchmark yields 99–100% interception on indirect injection, confirming that the TC1/TC2 capability is robust and not overfit to our synthetic data.
RQ2 (Security–latency trade-off). A two-stage cascade (Qwen3.5-9BGPT-4o-mini) achieves 91.9–92.6% IR with 1.9–6.7% FPR. For fully data-sovereign deployments, a local-only cascade (Qwen3.5-9BQwen2.5-14B) achieves 94.7–95.6% IR with 6.0–9.7% FPR (), confirming that a viable local alternative exists without cloud dependency. Qwen3.5-4B (3.4 GB, 490 ms) offers a viable edge-deployment option with 94.3–95.8% IR, though its elevated FPR (27.5–29.2%) limits suitability to risk-tolerant or human-in-the-loop settings, while Qwen2.5-14B approaches cloud-level precision on Chinese (FPR 9.7% vs. GPT-4o-mini’s 3.2%) at local-deployment cost.

### 10 Mar 2026 v2

@@ 7.4.6. Quantifying the GPT-4o-mini Circularity Bias
Since GPT-4o-mini generated the benchmark and also serves as one of the evaluated judges, its interception rates may be inflated by familiarity with its own output distribution. We bound this bias by comparing GPT-4o-mini’s performance against Qwen3.5-9B, an independently developed model from a different vendor (Alibaba vs. OpenAI) with a distinct training pipeline. On the Chinese dataset, Qwen3.5-9B achieves 96.5% IR versus GPT-4o-mini’s 93.0%; on English, 95.0% versus 95.4%. That an independent model achieves equal or higher interception suggests that the benchmark captures genuine attack patterns rather than GPT-4o-mini-specific artifacts. Moreover, GPT-4o-mini’s lower interception on Chinese (93.0% vs. 96.5%) contradicts the circularity hypothesis: if same-source bias were dominant, GPT-4o-mini should consistently outperform, which it does not. On IR, GPT-4o-mini exceeds Qwen3.5-9B only on English (95.4% vs. 95.0%, +0.4 pp) while underperforming on Chinese (93.0% vs. 96.5%, pp); the inconsistent direction across languages indicates that any same-source IR inflation is small relative to the inter-model variance. On FPR, GPT-4o-mini’s lower rate (3.2–12.3% vs. 23.7–37.7%) could partly reflect familiarity with its own benign outputs; this effect cannot be isolated without regenerating the benchmark with an independent generator. We note that this analysis bounds output-distribution familiarity on IR; a deeper structural bias—whereby GPT-4o-mini may also recognize its own attack-pattern preferences embedded in the generation templates—is harder to bound and remains a limitation. An analogous bias on FPR, whereby GPT-4o-mini may more reliably classify its own benign outputs as safe, cannot be excluded; bounding either effect would require regenerating the benchmark with an independent model (e.g., Claude or Gemini).
@@ 7.6. Generalization to External Benchmark
Table 10 shows that both judges achieve near-perfect interception on InjecAgent (GPT-4o-mini: 100% [95% CI: 96.4–100%]; Qwen3.5-9B: 99% [95% CI: 94.6–100%]), matching or exceeding their performance on our synthetic indirect injection subset. This provides evidence that the intent verification approach generalizes to independently constructed attack scenarios and is not overfit to our benchmark’s generation process. Note that these results measure judge-level interception on isolated tool calls, not end-to-end agent compromise reduction; validating LGA’s impact on full-pipeline attack success rates remains future work.
@@ 8. Conclusion
The experimental findings in this paper support the architectural argument that governance boundaries—sandbox isolation, intent verification, zero-trust protocols, audit logging—are necessary complements to model capability improvements in autonomous agent deployments. We formalize three security threat classes in multi-agent systems—Agency Abuse, RAG Data Poisoning, and Malicious Skill Plugins—and propose LGA, a four-layer governance architecture with empirical evaluation across all three.
RQ1 (Intent verification effectiveness). Across four local LLM judges, two NLI baselines, and one cloud judge evaluated on a 1,081-sample bilingual benchmark, all five LLM-based judges achieve at least 93% overall interception on TC1/TC2, while lightweight NLI baselines remain below 10%. Qwen2.5-14B achieves the best local balance (98.2–98.5% IR, 9.7–20.1% FPR), approaching GPT-4o-mini (93.0–95.4% IR, 3.2–12.3% FPR). On TC3 (malicious skill plugins), interception among judges with meaningful precision–recall balance drops to 75–94%; judges at the high end of the raw range (e.g., Llama-3.1-8B ZH at 99.2%) achieve this via conservative blocking bias (87.5% FPR on benign plugin calls; see Sect. 7.4.3), suggesting that permission-boundary violations are harder to detect from intent alignment alone (; TC3 results warrant larger-scale validation), motivating complementary enforcement at Layers 1 and 3 (sandbox isolation and schema validation). Despite this TC3 limitation, generalization to the external InjecAgent benchmark yields 99–100% interception on indirect injection, providing preliminary evidence that the TC1/TC2 interception capability generalizes beyond our synthetic benchmark, subject to the evaluation protocol limitations noted above.
RQ2 (Security–latency trade-off). A two-stage cascade (Qwen3.5-9BGPT-4o-mini) achieves 91.9–92.6% IR with 1.9–6.7% FPR. For fully data-sovereign deployments, a local-only cascade (Qwen3.5-9BQwen2.5-14B) achieves 94.7–95.6% IR with 6.0–9.7% FPR (), indicating that a viable local alternative exists without cloud dependency. Qwen3.5-4B (3.4 GB, 490 ms) offers a viable edge-deployment option with 94.3–95.8% IR, though its elevated FPR (27.5–29.2%) limits suitability to risk-tolerant or human-in-the-loop settings, while Qwen2.5-14B approaches cloud-level precision on Chinese (FPR 9.7% vs. GPT-4o-mini’s 3.2%) at local-deployment cost.